OpenStudy (anonymous):

1/2 * x + 2 = 10 1/2 is a fraction im not dividing

OpenStudy (hlambach):

I assume you're trying to find x?

OpenStudy (anonymous):

yes

OpenStudy (hlambach):

Okay. So start with the 2, you need to get it to the other side.

OpenStudy (anonymous):

first subtract x to do inverse

OpenStudy (anonymous):

not x imeant 2

OpenStudy (anonymous):

now u have 1/2x=8 what do u have to do to make 1/2x to 1x

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

well 2 1/2 make a whole right?

OpenStudy (hlambach):

1/2x + 2 = 10 -2 -2 What do you get?

OpenStudy (anonymous):

where did you get the other 2

OpenStudy (anonymous):

so multiply by 2 to 8

OpenStudy (anonymous):

to get x=?

OpenStudy (anonymous):

i got 16 but it was wrong

OpenStudy (hlambach):

The opposite of 1/2 is 2, which is why you need to multiply 8 by 2.

OpenStudy (anonymous):

and where did you get the other 2??

OpenStudy (anonymous):

the other 2 is to show inverse to both sides

OpenStudy (anonymous):

but the 1/2 is not division its a fraction

OpenStudy (anonymous):

we didnt divide

OpenStudy (anonymous):

flip 1/2 to 2/1

OpenStudy (anonymous):

2/1=2

OpenStudy (hlambach):

You're trying to get 1/2 to the other side. So we're doing the opposite, just like we did with the +2 part in the equation. Instead of ADDING 2, we SUBTRACTED 2. Now to get x by itself, you need to multiply 1/2 by it's opposite. We are multiplying because fractions are really just division problems. So we multiply each side by 2. Got it?

OpenStudy (anonymous):

but this one isnt division

OpenStudy (hlambach):

Like I said, fractions are just division problems. What you're doing to the 1 is DIVIDING it by 2.

OpenStudy (hlambach):

1 divided by 2 = 1/2

OpenStudy (hlambach):

Okay, let's try it a "different" way. You are multiplying x by 1/2, right? So do the opposite, DIVIDE both sides by 1/2.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

but i tried that and it was wrong

OpenStudy (hlambach):

It's right.

OpenStudy (hlambach):

Plug in 16 for x and you'll see it's right. 1/2(16) + 2 = 10

OpenStudy (hlambach):

1/2 * 16 = 8 8 + 2 = 10

OpenStudy (anonymous):

but its said it was -16 after i got it wrong

OpenStudy (hlambach):

Unless the original equation is -1/2x +2 = 10 You are right.

OpenStudy (anonymous):

ok

OpenStudy (hlambach):

If I helped, consider giving me a medal, please. :)

OpenStudy (hlambach):

Thanks!

OpenStudy (anonymous):

your welcome